When a fair die is rolled, it has equal chance to show one of six faces, labelled 1 through 6. If it is rolled 10 times, what is the probability that there are at least three 6's?
Answer #1

TOPIC:Binomial distribution.

=> Define, = P (of getting a 6 in a soll) wool) = There are total 6 Now, n = No. of times, the die was rolled. / possible out= 1- [P(x=0) +P(x =D +P(x=2) --[(0)4+)* (33 + (1999 +(2) (+)?( By the bol of x 0 2248 (in four decimale. (Aus)